# Spectra and pseudospectra in the evaluation of material stability

### v2

@@ 3.2 External power and invariance requirement: consequences
A subset $\mathfrak{b}$ of $\mathcal{B}$ is said to be a part of $\mathcal{B}$ when it has non-vanishing volume and the same geometrical regularity required of $\mathcal{B}$.
External actions over $\mathfrak{b}$ are associated with $u$ and $\nu$; they are defined by the power that they perform in the time rates of the two fields considered. Such actions are subdivided, as usual, into bulk and contact families: those associated with $u$ are standard bulk forces and boundary traction, while those performing power in $\dot{\nu}$ are microstructural actions. The external power $\mathcal{P}^{ext}_{\mathfrak{b}}$ that they perform on any time rate of change in the body morphology is defined in Lagrangian representation to be
@@
for any choice of $\mathfrak{b}$, $\mathfrak{c}$, and $q$. It means that we consider $\mathcal{P}_{\mathfrak{b}}^{ext}$ to be objective. The arbitrariness of $\mathfrak{c}$ implies the standard integral balance of forces
@@
Equation (1) gives standard results. Precisely, if $\left|b^{{\ddagger}}\right|$ is bounded over $\mathcal{B}$ and $\mathfrak{t}_{\partial}$ depends continuously on $x$, Cauchy’s theorem on fluxes implies that $\mathfrak{t}_{\partial}$ depends on $\mathfrak{\partial\mathfrak{b}}$ only through the normal $\mathbf{n}$ to it in all points where $\mathbf{n}$ is well-defined; $\mathfrak{t}_{\partial}$ is also such that $\mathfrak{t}_{\partial}=\mathfrak{t}:=\tilde{\mathfrak{t}}\left(x,t,\mathbf{n}\right)=-\tilde{\mathfrak{t}}\left(x,t,-\mathbf{n}\right)$, and, as a function of $\mathbf{n}$, $\tilde{\mathfrak{t}}$ is homogeneous and additive, namely $\tilde{\mathfrak{t}}\left(x,t,\mathbf{n}\right)=P\left(x,t,\right)\mathbf{n}\left(x\right)$, with $P$ the first Piola-Kirchhoff stress.
Less standard are the consequences of the integral balance (2). Precisely, since $\mathcal{B}$ is bounded, we can choose the arbitrary point $y_{0}$ in a way such that the boundedness of $|b^{{\ddagger}}|$ implies the one of $|(y-y_{0})\times b^{{\ddagger}}|$. If $\left|\mathcal{A}^{\top}\beta^{{\ddagger}}\right|$ is also bounded and $\tau_{\partial}$ depends continuously on $x$, by exploiting once again Cauchy’s theorem, we realize that $\tau_{\partial}$ depends on $\mathfrak{\partial\mathfrak{b}}$ only through the normal $\mathbf{n}$; also, $\mathcal{A}^{\top}\left(\tilde{\tau}\left(x,t,\mathbf{n}\right)+\tilde{\tau}\left(x,t,-\mathbf{n}\right)\right)=0$ and, as a function of $\mathbf{n}$, $\tilde{\tau}$ is homogeneous and additive: $\tilde{\tau}\left(x,t,\mathbf{n}\right)=\mathcal{S}\left(x,t\right)\mathbf{n}\left(x\right)$. The second-rank tensor $\mathcal{S}$ is named a microstress.
Due to the arbitrariness of $\mathfrak{b}$, if both stress fields are in $C^{1}\left(\mathcal{B}\right)\cap C\left(\bar{\mathcal{B}}\right)$ and the bulk actions $x\longmapsto b$, $x\longmapsto\beta^{{\ddagger}}$ are continuous over $\mathcal{B}$, from (1) we get the standard point-wise balance of forces
@@
where $\mathrm{skw}(\cdot)$ extract the skew-symmetric part of its argument; moreover,
@@
In the case of quasicrystals that we treat here with details, $\tilde{\nu}$ coincides with the so-called phason field; for such a material the possibility of peculiar inertia effects attributed to $\tilde{\nu}$ is questionable since only three sound-like branches are apparently recorded in experiments [33]. Consequently, even if only for later purposes, we leave apart the microstructural inertia and set $\kappa(\nu,\dot{\nu})=0$, so $\beta^{in}=0$ to within a powerless term that we also discard (for its presence in quasicrystals see [25] and [2]). The gross kinetic energy remains; the arbitrariness of $\dot{y}$ implies by continuity the identification $b^{in}=-\rho\ddot{y}=-\rho\ddot{u}$.
@@ 4 Constitutive structures
where $\tilde{\varsigma}(\dots)$ is a positive-valued state function, here chosen to be a scalar $\varsigma>0$ only for the sake of simplicity. With $\nu$ referring to a length scale $\lambda$, both $\mathcal{S}$ and $\mathbf{z}$ are thus of order $\lambda$. Consequently, from equation (4), we obtain that $PF^{\top}$ is symmetric to within terms of the order $\lambda^{2}$.
$P$, $\mathcal{S}$, and $\mathbf{z}$ are values of fields defined over $\mathcal{B}$ in terms of space variables. They have counterparts defined over $\mathcal{B}_{c}:=\tilde{y}(\mathcal{B},t)$ and given by
$\sigma=\frac{1}{\mathrm{det}F}\frac{\partial\psi}{\partial F}F^{\top}\;,\quad\mathbf{z}_{c}=\frac{1}{\mathrm{det}F}\left(\frac{\partial\psi}{\partial\nu}+\tilde{\varsigma}(\dots)\dot{\nu}\right)\;,\quad\mathcal{S}_{c}=\frac{1}{\mathrm{det}F}\frac{\partial\psi}{\partial N}F^{\top},$
@@
In small strain regime, that is, we recall, when $|\nabla u|\ll 1$, we have the approximations
$\sigma\approx P\;,\quad\mathbf{z}_{c}\approx\mathbf{z}\;,\quad\mathcal{S}_{c}\approx\mathcal{S}\;.$
@@
$+k_{3}\left(\mathrm{sym}\nabla u\cdot I\right)\left(N\cdot I\right)+k_{3}^{\prime}\,\mathrm{sym}N\cdot\mathrm{sym}\nabla u+\frac{1}{2}k_{0}\left|\nu\right|^{2}$
@@
$P\approx\boldsymbol{\sigma}=\lambda\left(\mathrm{tr}\,(\mathrm{sym}\nabla u)\right)I+2\mu\mathrm{sym}\nabla u+k_{3}\left(\mathrm{tr}\nabla\nu\right)I+k_{3}^{\prime}\mathrm{sym}\nabla\nu+\epsilon\nabla\dot{u}\;,$
@@
$c\dot{\nu}+k_{0}\nu=\chi\Delta u+\alpha\nabla\mathrm{div}\,u+\zeta\Delta\nu+\gamma\nabla\mathrm{div}\,\nu\;.$ (7)
@@ 5 Waves with time-dependent amplitude
$u(x,t)=\mathbf{u}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\quad\text{and}\quad\nu(x,t)=\mathbf{\nu}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\;,$ (8)
where, we repeat, $\mathbf{k}=k\mathbf{n}$ is a constant vector, the wave amplitudes $\mathbf{u},\mathbf{\nu}\in\mathbb{R}^{3}$ depend on time, and $\mathbf{x}=x-x_{0}$, with $x_{0}$ an arbitrary point.
We consider various possible choices for the phason self-action in order to evaluate the effects of the action itself.
@@ 5.1 In the absence of a self-action
$\sigma(A)=\bigcup_{\lambda\in\sigma(K)}\pm\sqrt{\lambda}\;.$
Hence, the eigenvalues of $A$ are either real or purely imaginary. In particular the eigenvalues of $A$ are real, both positive and negative, when $\mathbb{K}$ is positive definite; imaginary when $\mathbb{K}$ is negative definite and real and imaginary when $\mathbb{K}$ is indefinite. Moreover, the spectrum of $A$ is stable under (symmetric) perturbations in $\mathbb{K}$ thanks to Weyl’s Theorem [16], and thus the eigenvalues provide reliable information on the stability of the system.
@@ 5.3 A purely dissipative self-action
Assume that $\mathbf{z}$ is endowed only with a dissipative component driving diffusion, namely it is of the form $\mathbf{z}=\varsigma\dot{\nu}$, with $\varsigma$ a positive constant. From the balance equations we obtain
@@
$\rho\ddot{\mathbf{u}}=\mathbb{K}_{1}\mathbf{u}+\frac{1}{\rho}\mathbb{K}_{2}\mathbf{\nu}\;,\qquad\dot{\mathbf{\nu}}=\frac{1}{\varsigma}\mathbb{K}_{2}\mathbf{u}+\mathbb{K}_{3}\mathbf{\nu}\;,$ (21)
@@ 5.4 Simultaneous presence of conservative and dissipative components in the self-action
$\mathbb{K}_{3}:=-k^{2}\bigg{(}\frac{\zeta+\frac{k_{0}}{k^{2}}}{c}I+\frac{\gamma}{c}(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\;.$ (31)
@@ 6.1 Reasons for calling upon the pseudospectrum
The analysis along these lines is straightforward when $A$ is a normal matrix. The eigenvalues indicate characteristic times of exponentially relaxing modes whose linear combination is a solution to the considered system. When small perturbations are accounted for, small is the variation of the $A$ spectrum. At variance, when $A$ is non-normal, as in cases we tackle here, small perturbations can have significant effects that may induce instability. More specifically, for matrices far from being normal the standard spectral analysis shows possible failures in the evaluation of resonance, asymptotic behavior, diagonalization, geometric characterization [36]. Moreover, little or no information on the transient behavior and rate of convergence to steady-state can be gleaned in the highly non-normal case from the eigenvalues alone [1].
@@ 6.2 Definitions and a few properties
Given $A\in\mathbb{C}^{n\times n}$, take an arbitrary $\varepsilon>0$. The $\varepsilon$-pseudospectrum of $A$, indicated by $\sigma_{\varepsilon}(A)$, is the set of $\zeta\in\mathbb{C}$ such that
$\|(\zeta I-A)^{-1}\|_{2}>\frac{1}{\varepsilon}\;.$ (32)
Here we use the convention $\|(\zeta I-A)^{-1}\|_{2}=\infty$ if $\zeta I-A$ is singular.
@@
$\|(\zeta I-A)\mathbf{v}\|_{2}<\epsilon$ (34)
@@
The idea – we repeat – is thus to compute the spectra of matrices differing from $A$ by a perturbation given by matrices in a ball of radius $\varepsilon$, evaluating in this way how far the spectrum of $A$ can go under perturbations.
Assume that $A$ is diagonalizable and is not necessarily normal. Indicate by $V$ the matrix of eigenvectors $\mathbf{v}$ of $A$, namely
$\kappa_{2}(V):=\|V\|_{2}\|V^{-1}\|_{2}=\frac{s_{\max}(V)}{s_{\min}(V)}\;,$ (35)
where $s_{\max}(V)$ and $s_{\min}(V)$ are the maximum and minimim singular values of $V$, is called the spectral condition number of $V$. In general, $1\leq\kappa_{2}(V)<\infty$; specifically, if $A$ is a normal matrix, we have $\kappa_{2}(V)=1$. It is known that if $A$ is far from normal, then $\kappa_{2}(V)$ will be large. While the converse is not true in general, an ill-conditioned $V$ often signals that at least one of the eigenvalues of $A$ will be sensitive to small perturbations. Hence, one can use the condition number of $V$, $\kappa_{2}(V)\gg 1$, as a measure of non-normality, albeit with caution.
Another tool for detecting possible eigenvalue instabilities is the computation of the $\varepsilon$-pseudospectrum for suitable values of $\varepsilon$. Given $\varepsilon>0$, the pseudospectrum of $A\in\mathbb{C}^{n\times n}$, namely $\sigma_{\varepsilon}(A)$, referred to the Euclidan norm, satisfies what follows (see [36] for the pertinent proofs):
@@
    $\forall c\in\mathbb{C}$, $\sigma_{\varepsilon}(A+cI)=c+\sigma_{\varepsilon}(A)$.
@@
    $\forall c\in\mathbb{C}$, $c\not=0$, we get
    $\sigma_{|c|\varepsilon}(cA)=c\sigma_{\varepsilon}(A)\;.$
    In particular, for $c=2$ we have
@@
Since the matrices involved in our analysis are real, it seems more appropriate to perturb the system matrix $A\in\mathbb{R}^{N\times N}$ with real perturbation matrices $E\in\mathbb{R}^{N\times N}$.
Definition 4
The real structured $\varepsilon$-pseudospectrum of a matrix $A\in\mathbb{R}^{N\times N}$ is defined as
$\sigma_{\varepsilon}^{\mathbb{R}}(A)=\bigcup_{\begin{subarray}{c}E\in\mathbb{R}^{N\times N}\\ \|E\|<\varepsilon\end{subarray}}\sigma(A+E)\;.$ (36)
The real $\varepsilon$-pseudospectrum is of course a subset of the complex $\varepsilon$-pseudospectrum, $\sigma_{\varepsilon}^{\mathbb{R}}(A)\subset\sigma_{\varepsilon}(A)$.

### v3

@@ 3.2 External power and invariance requirement: consequences
A subset $\mathfrak{b}$ of $\mathfrak{B}$ is said to be a part of $\mathfrak{B}$ when it has non-vanishing volume and the same geometrical regularity required of $\mathfrak{B}$.
External actions over $\mathfrak{b}$ are associated with $u$ and $\nu$; they are defined by the power that they perform in the time rates of the two fields considered. Such actions are subdivided, as usual, into bulk and contact families: those associated with $u$ are the standard bulk force and boundary traction, while those performing power in $\dot{\nu}$ are the microstructural actions. The external power $\mathcal{P}^{ext}_{\mathfrak{b}}$ that they perform on any time rate of change in the body morphology is defined in Lagrangian representation to be
@@
for any choice of $\mathfrak{b}$, $c$, and $q$. It means that we consider $\mathcal{P}_{\mathfrak{b}}^{ext}$ to be objective. The arbitrariness of $c$ implies the standard integral balance of forces
@@
Equation (1) gives standard results. Precisely, if $|b^{{\ddagger}}|$ is bounded over $\mathfrak{B}$ and $\mathfrak{t}_{\partial}$ depends continuously on $x$, Cauchy’s theorem on fluxes implies that $\mathfrak{t}_{\partial}$ depends on $\mathfrak{\partial\mathfrak{b}}$ only through the normal $\mathbf{n}$ to it in all points where $\mathbf{n}$ is well-defined; $\mathfrak{t}_{\partial}$ is also such that $\mathfrak{t}_{\partial}=\mathfrak{t}:=\tilde{\mathfrak{t}}\left(x,t,\mathbf{n}\right)=-\tilde{\mathfrak{t}}\left(x,t,-\mathbf{n}\right)$, and, as a function of $\mathbf{n}$, $\tilde{\mathfrak{t}}$ is homogeneous and additive, namely $\tilde{\mathfrak{t}}\left(x,t,\mathbf{n}\right)=P\left(x,t,\right)\mathbf{n}\left(x\right)$, with $P$ the first Piola-Kirchhoff stress.
Less standard are the consequences of the integral balance (2). Precisely, since $\mathfrak{B}$ is bounded, we can choose the arbitrary point $y_{0}$ in a way such that the boundedness of $|b^{{\ddagger}}|$ implies the one of $|(y-y_{0})\times b^{{\ddagger}}|$. If $|\mathcal{A}^{\top}\beta^{{\ddagger}}|$ is also bounded and $\tau_{\partial}$ depends continuously on $x$, by exploiting once again Cauchy’s theorem, we realize that $\tau_{\partial}$ depends on $\mathfrak{\partial\mathfrak{b}}$ only through the normal $\mathbf{n}$; also, $\mathcal{A}^{\top}\left(\tilde{\tau}\left(x,t,\mathbf{n}\right)+\tilde{\tau}\left(x,t,-\mathbf{n}\right)\right)=0$ and, as a function of $\mathbf{n}$, $\tilde{\tau}$ is homogeneous and additive: $\tilde{\tau}\left(x,t,\mathbf{n}\right)=\mathcal{S}\left(x,t\right)\mathbf{n}\left(x\right)$. The second-rank tensor $\mathcal{S}$ is named a microstress.
Due to the arbitrariness of $\mathfrak{b}$, if both stress fields are in $C^{1}\left(\mathfrak{B}\right)\cap C\left(\bar{\mathfrak{B}}\right)$ and the bulk actions $x\longmapsto b$, $x\longmapsto\beta^{{\ddagger}}$ are continuous over $\mathfrak{B}$, from (1) we get the standard point-wise balance of forces
@@
where $\mathrm{skw}(\cdot)$ extracts the skew-symmetric part of its argument; moreover,
@@
In the case of quasicrystals that we treat here with details, $\tilde{\nu}$ coincides with the so-called phason field; for such a material the possibility of peculiar inertia effects attributed to $\tilde{\nu}$ is questionable since only three sound-like branches are apparently recorded in experiments [45]. Consequently, even if only for later purposes, we leave apart the microstructural inertia and set $\kappa(\nu,\dot{\nu})=0$, so $\beta^{in}=0$ to within a powerless term that we also discard (for its presence in quasicrystals see [34] and [2]). The gross kinetic energy remains; the arbitrariness of $\dot{y}$ implies by continuity the identification $b^{in}=-\rho\ddot{y}=-\rho\ddot{u}$ to within a powerless term that we also discard assuming that we are using an inertial frame of reference (in turn we can say that a frame of reference is inertial when the above identification of $b^{in}$ is exact).
@@ 4 Constitutive structures
where $\tilde{\varsigma}(\dots)$ is a positive-valued state function, here chosen to be a scalar $\varsigma>0$ only for the sake of simplicity. With $\nu$ referring to a length scale $\lambda$ (as already recalled in Section 2), both $\mathcal{S}$ and $\mathbf{z}$ are thus of order $\lambda$. Consequently, from equation (4), we obtain that $PF^{\top}$ is symmetric to within terms of the order $\lambda^{2}$.
$P$, $\mathcal{S}$, and $\mathbf{z}$ are values of fields defined over $\mathfrak{B}$ in terms of space variables. They have counterparts defined over $\mathfrak{B}_{c}:=\tilde{y}(\mathfrak{B},t)$ and given by
$\boldsymbol{\sigma}=\frac{1}{\mathrm{det}F}\frac{\partial\psi}{\partial F}F^{\top}\;,\quad\mathbf{z}_{c}=\frac{1}{\mathrm{det}F}\left(\frac{\partial\psi}{\partial\nu}+\tilde{\varsigma}(\dots)\dot{\nu}\right)\;,\quad\mathcal{S}_{c}=\frac{1}{\mathrm{det}F}\frac{\partial\psi}{\partial N}F^{\top},$
@@
In small strain regime, that is, we recall, when $\|\nabla u\|_{F}\ll 1$, we have the approximations
$\boldsymbol{\sigma}\approx P\;,\quad\mathbf{z}_{c}\approx\mathbf{z}\;,\quad\mathcal{S}_{c}\approx\mathcal{S}\;.$
@@
$+k_{3}\left(\mathrm{sym}\nabla u\cdot I\right)\left(N\cdot I\right)+k_{3}^{\prime}\,\mathrm{sym}N\cdot\mathrm{sym}\nabla u+\frac{1}{2}k_{0}\left\|\nu\right\|^{2}$
@@
$P\approx\boldsymbol{\sigma}=\lambda\left(\mathrm{tr}\,(\mathrm{sym}\nabla u)\right)I+2\mu\mathrm{sym}\nabla u+k_{3}\left(\mathrm{tr}\nabla\nu\right)I+k_{3}^{\prime}\mathrm{sym}\nabla\nu\;,$
@@
$\varsigma\dot{\nu}+k_{0}\nu=\chi\Delta u+\alpha\nabla\mathrm{div}\,u+\zeta\Delta\nu+\gamma\nabla\mathrm{div}\,\nu\;.$ (7)
@@ 5 Waves with time-dependent amplitude
$u(x,t)=\mathrm{Re}\{\mathbf{u}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\}\quad\text{and}\quad\nu(x,t)=\mathrm{Re}\{\mathbf{\nu}(t)e^{i\mathbf{k}\cdot\mathbf{x}}\}\;,$ (8)
where, we repeat, $\mathbf{k}=k\mathbf{n}$ is a constant vector, the wave amplitudes $\mathbf{u},\mathbf{\nu}\in\mathbb{C}^{3}$ depend on time, and $\mathbf{x}=x-x_{0}$, with $x_{0}$ an arbitrary point.
We consider various possible choices for the phason self-action in order to evaluate relevant effects.
@@ 5.1 In the absence of a self-action
$\sigma(A)=\bigcup_{\lambda\in\sigma(\mathbb{K})}\pm\sqrt{\lambda}\;.$
Hence, the eigenvalues of $A$ are either real or purely imaginary. In particular the eigenvalues of $A$ are real, both positive and negative, when $\mathbb{K}$ is positive definite; imaginary when $\mathbb{K}$ is negative definite and real and imaginary when $\mathbb{K}$ is indefinite. Moreover, the spectrum of $A^{2}$ is stable under (symmetric) perturbations in $\mathbb{K}$ thanks to Weyl’s theorem [21], and thus the eigenvalues provide reliable information on the stability of the system.
@@ 5.3 A purely dissipative self-action
Assume that $\mathbf{z}$ is endowed only with a dissipative component driving diffusion, namely it is of the form $\mathbf{z}=\varsigma\dot{\nu}$, with $\varsigma$ a positive constant, as chosen in Section 4. From the balance equations we obtain
@@
$\ddot{\mathbf{u}}=\mathbb{K}_{1}\mathbf{u}+\frac{1}{\rho}\mathbb{K}_{2}\mathbf{\nu}\;,\qquad\dot{\mathbf{\nu}}=\frac{1}{\varsigma}\mathbb{K}_{2}\mathbf{u}+\mathbb{K}_{3}\mathbf{\nu}\;,$ (21)
@@ 5.4 Simultaneous presence of conservative and dissipative components in the self-action
$\mathbb{K}_{3}:=-k^{2}\bigg{(}\frac{\zeta+\frac{k_{0}}{k^{2}}}{\varsigma}I+\frac{\gamma}{\varsigma}(\mathbf{n}\otimes\mathbf{n}\mathbf{)}\bigg{)}\;.$ (31)
@@ 6.1 Reasons for calling upon the pseudospectrum
The analysis along these lines is straightforward when $A$ is a normal matrix. The eigenvalues indicate characteristic times of exponentially relaxing modes whose linear combination is a solution to the considered system. When small perturbations are accounted for, small is the variation of the $A$ spectrum. At variance, when $A$ is non-normal, as in the cases we tackle here, small perturbations can have significant effects that may induce instability. More specifically, for matrices far from being normal the standard spectral analysis shows possible failures in the evaluation of resonance, asymptotic behavior, diagonalization, geometric characterization [49]. Moreover, little or no information on the transient behavior and rate of convergence to steady-state can be gleaned in the highly non-normal case from the eigenvalues alone [1].
@@
6.2 Scalar measures of departure from normality
Let $A\in\mathbb{C}^{n\times n}$ be a non-normal matrix and let $\mathfrak{K}$ be the set of $n\times n$ normal matrices. A measure of departure from normality is a real positive valued function on the space of matrices $\text{dep}:\mathbb{C}^{n\times n}\rightarrow\mathbb{R}^{+}$ such that $\text{dep}(A)=0$ if and only if $A\in\mathfrak{K}$.
The most natural measure of non-normality is given by the distance between $A$ and the set $\mathfrak{K}$, namely
${\text{dist}}(A,\mathfrak{K})=\inf_{N\in\mathfrak{K}}\|A-N\|\;,$ (32)
[…]
Assume that $A$ is diagonalizable and not necessarily normal. Indicate by $V$ a matrix of eigenvectors $\mathbf{v}$ of $A$, namely
$\kappa_{2}(V):=\|V\|_{2}\|V^{-1}\|_{2}=\frac{s_{\max}(V)}{s_{\min}(V)}\;,$ (33)
where $s_{\max}(V)$ and $s_{\min}(V)$ are the largest and smallest singular values of $V$, is called the spectral condition number of $V$. In general, $1\leq\kappa_{2}(V)<\infty$; specifically, if $A$ is a normal matrix, we have $\kappa_{2}(V)=1$. It is known that if $A$ is far from normal, then $\kappa_{2}(V)$ will be large. While the converse is not true in general, an ill-conditioned $V$ often signals that at least one of the eigenvalues of $A$ will be sensitive to small perturbations. Hence, one can use the spectral condition number of $V$, $\kappa_{2}(V)\gg 1$, as a measure of non-normality, albeit with caution.
If $A$ is normal, the commutator $[A,A^{*}]=AA^{*}-A^{*}A$ vanishes. Thus, one more natural measure of non-normality is given by the norm of the commutator between the matrix $A$ and its conjugate transpose $A^{\ast}$, divided by the square of the norm of $A$, to make the measure scale invariant, namely
$\text{dep}^{c}(A):=\frac{\|A^{\ast}A-AA^{\ast}\|_{2}}{\|A\|_{2}^{2}}\;.$ (34)
It has been proved (see [13]) that $0\leq\text{dep}^{c}(A)\leq\sqrt{2}$. Of course, $\text{dep}^{c}(A)=0$ if and only if $A$ is normal. Another measure of departure from normality, introduced by Henrici in [20], is based on the Schur decomposition. Let $A\in\mathbb{C}^{n\times n}$ be any matrix and let $A=U(D+T)U^{*}$ be the Schur decomposition of $A$, where $D=\text{diag}(\lambda_{1},\dots,\lambda_{n})$ is a diagonal matrix whose main diagonal is made up of the eigenvalues of $A$ and $T$ is a strictly upper triangular matrix. A classical result is that $T=0$ if and only if $A$ is normal. Therefore, Henrici’s departure from normality is defined as
$\text{dep}^{H}(A):=\inf_{A=U(D+T)U^{*}}\|T\|\;,$ (35)
where the infimum is taken over all possible Schur decompositions. However, working in the Frobenius norm this quantity becomes independent on the particular decomposition: the quantity
$\text{dep}^{H}_{F}(A)=\|T\|_{F}=\sqrt{\|A\|_{F}^{2}-\|D\|_{F}^{2}}$ (36)
is the same for any decomposition $A=U(D+T)U^{*}$ of $A$.
Using Henrici’s departure from normality it is possible to bound the distance of $A$ from the set of normal matrices $\mathfrak{K}$ both from above and below. Indeed, it holds that (see [14] and [25] for the proof)
$\frac{1}{\sqrt{n}}\text{dep}^{H}_{F}(A)\leq{\text{dist}}(A,\mathfrak{K})\leq\text{dep}^{H}_{F}(A)\;,$ (37)
where the distance is computed in the Frobenius norm. However, departure from normality does not imply necessarily eigenvalue instability under perturbations. The computation of the complex and of the real-structured $\varepsilon$-pseudospectra, recalled in the following section, for suitable values of $\varepsilon$ is a tool for detecting possible eigenvalue instabilities due to matrix non-normality.
@@ 6.3 Definitions and a few properties
Given $A\in\mathbb{C}^{n\times n}$, take an arbitrary $\varepsilon>0$. The $\varepsilon$-pseudospectrum of $A$, indicated by $\sigma_{\varepsilon}(A)$, is the set of $\hat{\zeta}\in\mathbb{C}$ such that
$\|(\hat{\zeta}I-A)^{-1}\|_{2}>\frac{1}{\varepsilon}\;.$ (38)
Here we use the convention $\|(\hat{\zeta}I-A)^{-1}\|_{2}=\infty$ if $\hat{\zeta}I-A$ is singular.
@@
$\|(\hat{\zeta}I-A)\mathbf{v}\|_{2}<\epsilon$ (40)
@@
The idea—we repeat—is thus to compute the spectra of matrices differing from $A$ by a perturbation given by matrices in a ball of radius $\varepsilon$, evaluating in this way how far the spectrum of $A$ can go under perturbations.
Given $\varepsilon>0$, the pseudospectrum of $A\in\mathbb{C}^{n\times n}$, namely $\sigma_{\varepsilon}(A)$, referred to the Euclidan norm, satisfies what follows (see [49] for the pertinent proofs):
@@
    $\forall\mathsf{h}\in\mathbb{C}$, $\sigma_{\varepsilon}(A+\mathsf{h}I)=\mathsf{h}+\sigma_{\varepsilon}(A)$.
@@
    $\forall\mathsf{h}\in\mathbb{C}$, $\mathsf{h}\not=0$, we get
    $\sigma_{|\mathsf{h}|\varepsilon}(\mathsf{h}A)=\mathsf{h}\sigma_{\varepsilon}(A)\;.$
    In particular, for $\mathsf{h}=2$ we have